| <p><strong>09:44</strong></p>
<p>Run 26 started - Beam no target</p>
<p>Beam energy is being increased from 1.5 to 3 MeV</p>
<p>Horizontal beam position = 7.0 mm, vertical beam position = 0.0 mm.</p>
<p>Detectors are positioned fully OUT</p>
<p>Histograms and stats zero'd, monitoring code restarted.</p>
<p>Pressure in YR09 E target ~ 4.5e-11 mbar.</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p> </p>
<p><strong>10:48 (Attachements 1-4)</strong></p>
<p> </p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>The detectors currently fully out for tuning.</p>
<p> </p>
<p> </p>
<p><strong>11:57 (attachments 5-8)</strong></p>
<p> </p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>The detectors currently fully out for tuning.</p>
<p> </p>
<p> </p>
<p> </p>
<p> </p> |